5 -kg crate is on a plane inclined 30 degree above the horizontal . A force of 20 nt is used to move it upwards . After how many seconds will the crate attain a velocity of 2.78 m/sec if the coefficient of friction between the crate and the plane is 0.30?
1
Expert's answer
2020-12-17T07:26:13-0500
"v=at=\frac{(F-mg(\sin{30}+\mu \cos{30}))}{m}t\\
2.78=\frac{(20-5(9.8)(\sin{30}+0.3 \cos{30}))}{5}t\\t=-0.8\ s<0\ s"

Thus, the body cannot move upwards:


"F<mg(\sin{30}+\mu \cos{30})"